World class marathon runners are known to run that distance (26.2 miles) in an average of 144 minutes with a standard deviation of 15 minutes. If we sampled a group of world class runners from a particular race, find the probability of the following: **(use 4 decimal places)** a.) The probability that one runner chosen at random finishes the race in less than 138 minutes. b.) The probability that 10 runners chosen at random have an average finish time of less than 138 minutes. c.) The probability that 50 runners chosen at random have an average finish time of less than 138 minutes.
Answer #1

µ = 144

sd = 15

a) P(X<138)=P(\frac{X-\mu}{\sigma}<\frac{138-\mu}{\sigma})

                                =P(Z<\frac{138-144}{15})

                                = P(Z < -0.4)

                                = 0.3446

b) n = 10

P(\bar{X}<138)=P(\frac{\bar{X}-\mu}{\sigma/\sqrt{n}}<\frac{138-\mu}{\sigma/\sqrt{n}})

                            =P(Z<\frac{138-144}{15/\sqrt{10}})

                            = P(Z < -1.26)

                            = 0.1038

c) n = 50

P(\bar{X}<138)=P(\frac{\bar{X}-\mu}{\sigma/\sqrt{n}}<\frac{138-\mu}{\sigma/\sqrt{n}})

                            =P(Z<\frac{138-144}{15/\sqrt{50}})

                            = P(Z < -2.83)

                            = 0.0023